the problem says find the missing side of each triangle. Side c is the hypotenuse. Side a and b are the legs. Round your answer to the nearest tenth if necessary. so if b=12 mi, and c is 13 mi what would a equal?
Answer by mananth(16946)

a^2+12^2=13^2
a^2=13^2-12^2
a^2=169-144
a^2=25
a= 5